Hi! So i have a question about adding radicals. I was gone for this day so i am not sure how to do it.
add [(x+5)/(x-2)]+[(-10x-36)/(x^2+4x-12)]

Answer by edjones(8007)

[(x+5)/(x-2)]+[(-10x-36)/(x^2+4x-12)]
=(x+5)/(x+2) - (10x+36)/((x-2)(x+6))
=((x+5)(x+6))-(10x+36)/((x-2)(x+6))
=(x^2+11x+30-10x-36)/((x-2)(x+6))
=(x^2+x-6)/((x-2)(x+6))
=((x+3)(x-2))/((x-2)(x+6))
=(x+3)/(x+6)
.
Ed